FIND AN EQUATION FOR THE LINE WITH SLOPE 1/2 AND y-INTERCEPT 3.
Answer by uma(370)

Equation to a line with slope 'm' and y intercept 'c' is y = mx + c
Here m = 1/2 and c = 3
So the required equation is y = (1/2)x + 3
==> 2y = x + 6 [multiplying by 2 throughout]
==> 2y - x - 6 = 0

Good Luck!!!